1.      if a households income falls from R20 000 to R17 000 and its consumption falls from R18 000 to R15 000, then: 


1
Expert's answer
2021-09-10T11:19:48-0400

It's marginal propensity on saving tends to be the proportion or amount of income utilized in saving compared to consumption.

marginal propensity"=\frac{change in consumption}{change in income}"


"= \frac{8000-9500}{1000-12000}"


"=\frac{-1500}{-2000}"


= 0.750